A supplement of an angle is six more than three times the angle's complement. Write an equation to represent the problem and solve for the angle, its complement, and its supplement.

Answer by ankor@dixie-net.com(22740)

A supplement of an angle is six more than three times the angle's complement. Write an equation to represent the problem and solve for the angle, its complement, and its supplement.
:
Let x = the angle
then
(180-x) = it's supplement
and
(90-x) = it's complement
:
The equation for the given statement:
(180-x) = 3(90-x) + 6
180 - x = 270 - 3x + 6
-x + 3x = 276 - 180
2x = 96
x = 96/2
x = 48 degrees is the angle
then
180 - 48 = 132 degrees is the supplement
and
90 - 48 = 42 degrees is the complement
;
:
Check our solutions in the given statement:
"A supplement of an angle is six more than three times the angle's complement."
132 = 3(42) + 6
132 = 126 + 6; confirms our solutions